I would like to see how you set up this equation?
A soccer team held a car wash to raise money. They earned 6.50 for each car they washed. The team raised a total of 234.00 by washing cars. Which equation can be used to solve for n, the number of cars the team washed?
A 6.5= 234/n
B 6.5/n=234
c. 234 = n/6.5
d. n/234=6.5

Answer by gonzo(654)

equation would be:
6.5 * n = 234
---
this can be expressed as:
6.5 = 234/n
or
n = 234/6.5
---
looks like the answer would be A.
A is 6.5 = 234/n which matches the first form that the original equation ca be expressed as.
---